OpenStudy (anonymous):

Please solve !|-10 - x| = 6

OpenStudy (anonymous):

This is the same as the other one. I'll get you started by removing the absolute value. \[-10 - x = 6 ~\text{and} -10 -x = -6\]You should get two values of x.

OpenStudy (anonymous):

ooo

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[\left| -10-x\right|=6\] remove the absolute value 10+x=6 eliminate x x=6-10 x=-4